# Applying the maximum entropy principle to neural networks enhances multi-species distribution models

### v3

@@ Appendix A Supplementary theoretical aspects
A.1 Relationship between Poisson loss and Maxent loss
For this normalisation, $\lambda$ and $y$ at site $i$ for a given species $j$ are expressed as the normalised values $\tilde{\lambda}_{ij}$ and $\tilde{y}_{ij}$
$\tilde{\lambda}_{ij}=\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}$ (11)
$\tilde{y}_{ij}=\frac{y_{ij}}{\sum_{k=1}^{K}y_{kj}}$ (12)
Thus the Poisson loss in a context where intensities and the number of occurrences are normalised per site can be formulated as follow:
$\mathcal{L}(\tilde{\lambda},\tilde{y})$ $=\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}\left(\tilde{\lambda}_{ij}-\tilde{y}_{ij}\log\tilde{\lambda}_{ij}\right)$ (13)
$=\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}\left(\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}-\frac{y_{ij}}{\sum_{k=1}^{K}y_{kj}}\log\left(\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}\right)\right)$ (14)
$=\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}\left(\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}\right)-\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}y_{ij}\log\left(\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}\right)$ (15)
$=\frac{1}{K}-\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}\left(\frac{y_{ij}}{\sum_{k=1}^{K}y_{kj}}\right)\log\left(\frac{\lambda_{ij}}{\sum_{k=1}^{K}\lambda_{kj}}\right)$ (16)
$=\frac{1}{K}+\mathcal{L}_{\mathcal{H}}(\lambda,y)$ (17)
A.2 The case of Maxent
Maximum entropy (Maxent) is a widely used method in species distribution modelling (SDM) that estimates the distribution of a species on the basis of environmental variables. The objective of Maxent is to approximate the probability distribution $p(x)$ of species occurrences over a set of sites using PO data. The method maximises the entropy of the predicted distribution, subject to environmental constraints derived from occurrence data.
The Maxent approach can be seen as a special case of a Poisson Point Process (PPP). A PPP models the occurrence of species as events in space, with a function $\lambda(x)$ describing the intensity of occurrences at location $x$. If we denote by $X$ the set of all possible sites, the probability density of species occurrence can be written as follows:
@@ A.2 The case of Maxent
Maxent is equivalent to estimating the intensity $\lambda(x)$ using the environmental variables at each site $x$. In Maxent, $\lambda(x)$ is modelled as a log-linear function of features (e.g. environmental covariates):
$\lambda(x)=\exp\left(\beta_{0}+\sum_{i=1}^{p}\beta_{i}f_{i}(x)\right)$ (19)
where $f_{i}(x)$ are the environmental characteristics at location $x$, and $\beta_{i}$ are the parameters to be learned. The objective is to estimate $\beta_{i}$ by maximising the probability of observed occurrences. Given a set of occurrence points $S=\{x_{1},x_{2},\dots,x_{n}\}$, the loss function of Maxent is to maximise the likelihood under the assumption of a Poisson process. The complete Maxent loss can be written as follows:
$\mathcal{L}(\beta)=-\frac{1}{n}\sum_{x\in S}\log p(x)+\lambda\sum_{i=1}^{p}|\beta_{i}|$ (20)
where $p(x)$ is the probability of occurrence at site $x$, $\lambda$ is a regularisation parameter, and $|\beta_{i}|$ is the L1 regularisation term that promotes sparsity in the learned coefficients. In Maxent, the partition function $Z(\beta)$ is used to normalise the predicted distribution so that the sum of the probabilities is equal to 1. It is defined by the following integral:
@@
The predicted probability of occurrence $p(x)$ is then:
$p(x)=\frac{\exp\left(\beta_{0}+\sum_{i=1}^{p}\beta_{i}f_{i}(x)\right)}{Z(\beta)}$ (22)
A.3 Minimising the loss on each batch implies minimising the full loss
In DeepMaxent, the term batch is not to be understood in its classical sense in machine learning or statistics. This is because the formula of the loss for a batch is not simply the restriction of the full loss to the terms of that batch, due to the partition functions. Thus, unlike in a classical setting, using the SGD with this type of batches might not necessarily lead to an estimator that is good regarding the full loss. To address this, we show below that an estimator that minimises all the batch-wise losses actually minimises the full loss, thus justifying the use of the SGD.
Notations: Note for any $n\in\mathbb{N}^{\star}$, $\Delta^{n}=\{t_{0},...,t_{n}\in\mathbb{R}^{n+1}\/\sum_{i=0}^{n}t_{i}=1,\;\forall i,t_{i}\geq 0\}$ the $n$-probability simplex, i.e. the space of probability distributions on a set of $n+1$ elements. Without loss of generality, we consider a single species with occurrence (pseudo-)count $y_{k}$ in site $k$. $\forall k\in[1,K],y_{k}>0$, consistently with our implementation, where $10<\delta<<$ 1 is added to each raw occurrence count to avoid numerical problems. For any $B\subset[1,K]$, we note $y_{B}:=\{y_{k}\}_{k\in B}$, and also apply this notation to $\lambda_{B}$. We express the batch cross-entropy loss function of DeepMaxent with our notation in equation 23.
@@ A.3 Minimising the loss on each batch implies minimising the full loss
Now, assume that an estimator of the intensity $\lambda^{*}_{[1,K]}$ ($\forall k\in[1,K],\lambda^{*}_{k}>0$) minimises our loss for any batch of size $n$ ($1<n<K$). Formally, $\lambda^{*}$ fulfils the property $P(n)$: $\forall B\subset[1,K]$ such that $|B|=n,\;\{\lambda^{*}_{k}/\sum_{i\in B}\lambda^{*}_{i}\}_{k\in B}=\{y_{k}/\sum_{i\in B}y_{i}\}_{k\in B}$
As a preliminary corollary, we have that $P(n)\Rightarrow P^{\prime}(n)$: $\forall B_{1},B_{2}\subset[1,K]/|B_{1}|=|B_{2}|=n$ and $B_{1}\cap B_{2}\neq\emptyset,\;\frac{\sum_{i\in B_{1}}\lambda_{i}^{*}}{\sum_{i\in B_{2}}\lambda_{i}^{*}}=\frac{\sum_{i\in B_{1}}y_{i}}{\sum_{i\in B_{2}}y_{i}}$, because from $P(n)$ we get $\frac{\lambda_{k}^{*}}{y_{k}}=\frac{\sum_{i\in B_{1}}\lambda_{i}^{*}}{\sum_{i\in B_{1}}y_{i}}=\frac{\sum_{i\in B_{2}}\lambda_{i}^{*}}{\sum_{i\in B_{2}}y_{i}}$ which leads to the result.
Property: $P(n)\Rightarrow P(K)$, that is, if $\lambda^{*}$ fulfils $P(n)$, it minimises as well the full loss $\mathcal{L}(\lambda^{*}_{[1,K]},y_{[1,K]})$.
@@
$P(n+1)\Leftrightarrow\forall B\subset[1,K],l\in[1,K],\,k\notin B$, we have $\{\lambda^{*}_{k}/\sum_{i\in B\cup\{l\}}\lambda^{*}_{i}\}_{k\in B\cup\{l\}}=\{y_{k}/\sum_{i\in B\cup\{l\}}y_{i}\}_{k\in B\cup\{l\}}$
@@
$\frac{\lambda^{*}_{i}}{\sum_{k\in B\cup l}\lambda^{*}_{k}}$ $=\frac{\lambda^{*}_{i}}{\sum_{k\in B\cup l\backslash j}\lambda^{*}_{k}+\lambda^{*}_{j}}$ (24)
$=\frac{\lambda^{*}_{i}/\sum_{k\in B}\lambda^{*}_{k}}{\frac{\sum_{k\in B\cup l\backslash j}\lambda^{*}_{k}}{\sum_{k\in B}\lambda^{*}_{k}}+\frac{\lambda^{*}_{j}}{\sum_{k\in B}\lambda^{*}_{k}}}$
=$\underset{P(n),P^{\prime}(n)}{=}\frac{y_{i}/\sum_{k\in B}y_{k}}{\frac{\sum_{k\in B\cup l\backslash j}y_{k}}{\sum_{k\in B}y_{k}}+\frac{y_{j}}{\sum_{k\in B}y_{k}}}$
$=\frac{y_{i}}{\sum_{k\in B\cup l\backslash j}y_{k}+y_{j}}$
$=\frac{y_{i}}{\sum_{k\in B\cup l}y_{k}}$
@@
Appendix B Dataset description
Appendix C A suggestion for cross-validating DeepMaxent using PO data
In this study, model calibration was performed using presence-only (PO) data, whereas model evaluation used presence-absence (PA) data with a specific metric, the Area Under the Curve (AUC). This difference introduces a fundamental challenge in choosing the appropriate evaluation metric in the cross-validation step using PO data. PO and PA data are collected under different conditions with bias related to inhomogeneous sampling effort in the case of PO data. To fine-tune the DeepMaxent model’s hyperparameters, spatial blocking was employed for cross-validation, drawing on geographical data to optimise the model’s performance. The cross-validation was performed following the approach of Zbinden et al. [2024], Roberts et al. [2017]. In more detail, we performed a spatially stratified 10-fold cross-validation using a grid of 25 spatial blocks. Each fold used a unique subset of blocks for validation, distinct from previous folds, with blocks selected randomly yet balanced to ensure an even distribution of presence data across folds.
@@ Appendix C A suggestion for cross-validating DeepMaxent using PO data
Batch Size 250, 2500
@@
Weight Decay $10^{-3}$, $10^{-2}$ , $10^{-1}$
To achieve this, all combinations of hyperparameter values visible in Table 4 were evaluated. The evaluation criterion used was the average AUC across all validation sets. The optimal model was obtained with two hidden layers, a learning rate of 0.0002, a weight decay of $10^{-2}$ and a batch size of 250.
Table 5: All combinations of hyperparameter values and their cross-validation results on PO data.
Batch Size Hidden Layers Weight Decay AUC_CV
250 1 0.0001 0.632
250 1 0.001 0.631
250 1 0.01 0.649
250 2 0.0001 0.654
250 2 0.001 0.659
250 2 0.01 0.673
250 3 0.0001 0.619
250 3 0.001 0.624
250 3 0.01 0.635
2500 1 0.0001 0.634
2500 1 0.001 0.631
2500 1 0.01 0.632
2500 2 0.0001 0.659
2500 2 0.001 0.659
2500 2 0.01 0.660
2500 3 0.0001 0.617
2500 3 0.001 0.619
2500 3 0.01 0.634
To obtain optimal hyperparameter values, one of the main challenges is to find the right combination of hyperparameters during cross-validation while maintaining computational efficiency. Techniques such as subsampling the dataset or limiting the training process to a reduced number of epochs can help to reduce the computational time for each configuration.
Appendix D Supplementary results
D.1 Batch size effect on model performance
The impact of batch size on the average AUC values obtained by the DeepMaxent method in the different regions of the dataset is illustrated in Table 6.
Table 6: Impact of batch size on DeepMaxent performance across different regions. The table presents the accuracy scores for each region (AWT, CAN, NSW, NZ, SA, SWI) and the average (avg) performance across all regions.
Batch size regions

### v4

@@ Appendix A Supplementary theoretical aspects
A.1 The relationship between Poisson, DeepMaxent and Maxent loss functions
We show here that the global minimizer of DeepMaxent’s loss for a species probability distribution across sites $(\tilde{\lambda}_{ij})_{i\in 1,...,K}$ is equivalent to the one of the Poisson regression estimate, even though the Poisson regression adds the estimation of a multiplicative factor for the intensity captured by the parameter $b_{j}$. Then, using the equivalence between the log-linear Poisson regression and Maxent (Renner and Warton [2013]), we show that DeepMaxent is a non-linear generalization of Maxent, i.e. it reduces to Maxent for a log-linear model where $g_{\theta}(x_{i})=x_{i}$.
We consider a single species $j$ for simplicity and the site index $i\in\{1,...,K\}$.d
The occurrence count of $j$ at site $i$ is noted $y_{ij}\in\mathbb{N}$. We further note the total count $y_{j}^{+}=\sum_{i=1}^{K}y_{ij}>0$ and normalised count $\tilde{y}_{ij}=y_{ij}/y_{j}^{+}$.
The model predicted intensity in cell $i$ is $\lambda_{ij}=\lambda_{j}(x_{i})\geq 0$ with our main notations, i.e. some parameterized function of the site covariates, not necessarily log-linear. We also note $Z_{j}=\sum_{i=1}^{K}\lambda_{ij}$ and $\tilde{\lambda}_{ij}=\lambda_{ij}/Z_{j}$.
First, we’ll show the equivalence between the Poisson loss (negative log-likelihood) and DeepMaxent’s loss. This equivalence is strongly related to the Poisson trick (Lee et al. [2017]), but we write the proof below with our notations for transparency.
The Poisson loss (negative log-likelihood) for species $j$ is written in equation 8 below consistently with equation 1 (we simply removed the constant factor):
@@ A.1 The relationship between Poisson, DeepMaxent and Maxent loss functions
Consider now the particular case where $\lambda_{ij}$ is a log-linear function of the covariates $x_{i}$, i.e. $g_{\theta}(x_{i})=x_{i}$, in the Poisson loss above. This is the negative log-likelihood of the log-linear Poisson regression that is equivalent to Maxent, as shown by theorem 1 of Renner and Warton [2013]. Hence, the equivalence that we have shown above means that the DeepMaxent loss and estimate are equivalent to the ones of Maxent in this simple particular case. More broadly, DeepMaxent’s loss generalizes Maxent’s loss by allowing a more flexible model for the distribution across sites, and preserving the link to the Poisson loss.
A.2 A modified batched SGD algorithm for DeepMaxent: The empirical minimizer of all batch-wise losses minimizes DeepMaxent’s loss
In DeepMaxent, the term batch is not to be understood in its classical sense in machine learning or statistics. This is because the formula of the loss for a batch is not simply the restriction of the full loss to the terms of that batch: The partition function, used to normalise the intensity, is also reduced to the terms of the batch. Thus, unlike in a classical setting, using the SGD with this type of batches might not necessarily lead to an estimator that is good regarding the full loss. To address this, we show below that the empirical estimator that minimises all the batch-wise losses is the empirical estimator minimizing the full loss. However, this estimator exactly predicts the data, which, in general, is neither possible given a constrained model nor what we want. Even though the property doesn’t guarantee that the actual model parameters minimizing all batch-wise losses will actually minimize the full loss, it gives us the insight that they should be close, which justifies this training procedure.
Notations: Note for any $n\in\mathbb{N}^{\star}$, $\Delta^{n}=\{t_{0},...,t_{n}\in(\mathbb{R}^{+})^{n+1}\/\sum_{i=0}^{n}t_{i}=1,\;\forall i,t_{i}\geq 0\}$ the $n$-probability simplex, i.e. the space of probability distributions on a set of $n+1$ elements. Without loss of generality, we consider a single species with occurrence (pseudo-)count $y_{k}$ in site $k$. $\forall k\in\{1,\dots,K\},y_{k}>0$, consistently with our implementation, where $0<\epsilon<<1$ is added to each raw occurrence count to avoid numerical problems. For any $B\subset\{1,\dots,K\}$, we note $y_{B}:=\{y_{k}\}_{k\in B}$, and also apply this notation to $\lambda_{B}$. We express the batch cross-entropy loss function of DeepMaxent with our notation in equation 9.
@@
Now, assume that an estimator of the intensity $\lambda^{*}_{\{1,\dots,K\}}$ ($\forall k\in\{1,\dots,K\},\lambda^{*}_{k}>0$) minimises our loss for any mini-batch of size $n$ ($1<n<K$). Formally, $\lambda^{*}$ fulfils the property $P(n)$: $\forall B\subset\{1,\dots,K\}$ such that $|B|=n,\;\{\lambda^{*}_{k}/\sum_{i\in B}\lambda^{*}_{i}\}_{k\in B}=\{y_{k}/\sum_{i\in B}y_{i}\}_{k\in B}$
As a preliminary corollary, we have that $P(n)\Rightarrow P^{\prime}(n)$: $\forall B_{1},B_{2}\subset\{1,\dots,K\}/\penalty 50|B_{1}|=|B_{2}|=n$ and $B_{1}\cap B_{2}\neq\emptyset,\;\frac{\sum_{i\in B_{1}}\lambda_{i}^{*}}{\sum_{i\in B_{2}}\lambda_{i}^{*}}=\frac{\sum_{i\in B_{1}}y_{i}}{\sum_{i\in B_{2}}y_{i}}$, because from $P(n)$ we get $\frac{\lambda_{k}^{*}}{y_{k}}=\frac{\sum_{i\in B_{1}}\lambda_{i}^{*}}{\sum_{i\in B_{1}}y_{i}}=\frac{\sum_{i\in B_{2}}\lambda_{i}^{*}}{\sum_{i\in B_{2}}y_{i}}$ which leads to the result.
Property: $P(n)\Rightarrow P(K)$, that is, if $\lambda^{*}$ fulfils $P(n)$, it minimises as well the full loss $\mathcal{L}(\lambda^{*}_{\{1,\dots,K\}},y_{\{1,\dots,K\}})$.
@@
$P(n+1)\Leftrightarrow\forall B\subset\{1,\dots,K\},l\in\{1,\dots,K\},\,k\notin B$, we have $\{\lambda^{*}_{k}/\sum_{i\in B\cup\{l\}}\lambda^{*}_{i}\}_{k\in B\cup\{l\}}=\{y_{k}/\sum_{i\in B\cup\{l\}}y_{i}\}_{k\in B\cup\{l\}}$
@@
[…]
@@
Appendix B DeepMaxent architecture
B.1 Residual MLP architecture
Figure 5: The residual neural network to estimate the intensity $\lambda$ from variable input ($x_{0}$), where $P$ as is the variable number input, where $C$ is the number of hidden layer nodes, and NN denotes the number of species (or target categories). The illustrated case involves two hidden layers. In the special case where there is only one hidden layer, no residual addition is applied.
DeepMaxent is architecture-agnostic, and in this work we chose a $g$ based on a residual neural network architecture (see Figure 5), which takes the vector of environmental variables $\mathbf{x}$ as input. $\mathbf{x}$ is transformed by a sequence of hidden layers $g=g_{1}\circ g_{2}\circ\dots\circ g_{L}$, whose number $L$ and size are the main hyper-parameters. Residual connections between adjacent hidden layers were incorporated introducing shortcuts that allow the input of a layer to bypass the non-linear transformations and be directly added to the output ($\mathbf{x}_{l+1}=\mathbf{x}_{l}+g_{l}(\mathbf{x}_{l})$[He et al., 2016]. The output of the last hidden layer ($g(\mathbf{x})$), used as our latent feature vector, leads to the output layer, which is composed of each species intensity measure $\mathbf{\lambda}_{j}(\mathbf{x})$.
B.2 Convolutional DeepMaxent architecture (ResNet18-based)
In addition to the multilayer perceptron architecture described above, DeepMaxent can accommodate feature extractors based on convolutional neural networks when the environmental covariates have an inherent spatial, spectral or temporal structure. We explored an instantiation of $g_{\theta}$ based on a modified ResNet-18 backbone [He et al., 2016], adapted from [Picek et al., 2025] specifically for the Landsat-Geoplant use case. The input tensor, representing multi-channel layers, is first normalised and then passed through a ResNet-18 in which the initial convolutional layer is adapted to accept $C_{\text{in}}$ channels instead of the standard three. As large-scale pooling may remove fine-grained spatial information relevant to species distributions, the initial max-pooling operation is disabled. The convolutional blocks learn spatially structured latent features at multiple scales, producing a 1000-dimensional representation after the final global pooling stage. This representation is further normalised and transformed through two fully connected layers to obtain the final latent feature vector $g_{\theta}(\mathbf{x})$, which plays the same role as in the MLP-based version of DeepMaxent. This convolutional variant enables the model to exploit local spatial patterns, textures, or temporal gradients present in high-resolution environmental rasters, while preserving the multi-species output structure described previously.
Appendix C Cross-validating DeepMaxent using PO data
In this study, model calibration was performed using presence-only (PO) data, whereas model evaluation used presence-absence (PA) data with a specific metric, the Area Under the Curve (AUC). This difference introduces a fundamental challenge in choosing the appropriate evaluation metric in the cross-validation step using PO data. PO and PA data are collected under different conditions with bias related to inhomogeneous sampling effort in the case of PO data. To fine-tune the DeepMaxent model’s hyper-parameters, spatial blocking was employed for cross-validation, drawing on geographical data to optimise the model’s performance. The cross-validation was performed following the approach of Zbinden et al. [2024], Roberts et al. [2017]. In more detail, we performed a spatially stratified 10-fold cross-validation using a grid of 25 spatial blocks. Each fold used a unique subset of blocks for validation, distinct from previous folds, with blocks selected randomly yet balanced to ensure an even distribution of presence data across folds.
@@ Appendix C Cross-validating DeepMaxent using PO data
Learning rate 0.0002, 0.002, 0.02
Mini-batch size 250, 2500
@@
To achieve this, all combinations of hyper-parameter values visible in Table 5 were evaluated. The evaluation criterion used was the average AUC across all validation sets. The best model was obtained with two hidden layers, a learning rate of 0.0002 and a mini-batch size of 250.
For DeepMaxent’s loss, the best scores were achieved with a learning rate of 0.0002, two hidden layers, and a mini-batch size of 250. However, for the Poisson and BCE losses, cross-validation suggests a different optimal learning rate (lr = 0.002).
Table 6: All combinations of hyper-parameter values and their cross-validation results on PO data for each loss.
Learning rate Mini-batch Size Hidden Layers AUC
BCE CE DeepMaxent Poisson
0.0002 250 1 0.582 0.645 0.662 0.568
0.0002 250 2 0.611 0.654 0.668 0.591
0.0002 250 3 0.600 0.640 0.653 0.588
0.0002 2500 1 0.518 0.586 0.622 0.505
0.0002 2500 2 0.548 0.619 0.645 0.535
0.0002 2500 3 0.540 0.609 0.634 0.525
0.002 250 1 0.649 0.647 0.654 0.644
0.002 250 2 0.656 0.645 0.659 0.655
0.002 250 3 0.652 0.646 0.646 0.650
0.002 2500 1 0.597 0.654 0.661 0.578
0.002 2500 2 0.616 0.649 0.661 0.596
0.002 2500 3 0.602 0.646 0.648 0.585
0.02 250 1 0.644 0.634 0.633 0.643
0.02 250 2 0.645 0.627 0.631 0.643
0.02 250 3 0.643 0.621 0.640 0.633
0.02 2500 1 0.637 0.644 0.634 0.609
0.02 2500 2 0.626 0.624 0.650 0.604
0.02 2500 3 0.596 0.621 0.643 0.582
Table 7: Comparison of BCE and Poisson loss performances by region-averaged AUC and general averaged AUC over all regions.
Regions
AWT CAN NSW NZ SA SWI avg
Results from the literature
lr=0.002; mini-batch size = 250; hidden-layers=2
BCE (using TGB) 0.699 0.658 0.731 0.722 0.802 0.809 0.737
Poisson (using TGB) 0.709 0.677 0.740 0.729 0.803 0.819 0.746
lr=0.0002; mini-batch size = 250; hidden-layers=2
BCE (using TGB) 0.723 0.726 0.743 0.739 0.803 0.846 0.763
Poisson (using TGB) 0.712 0.727 0.732 0.731 0.800 0.846 0.758
Although cross-validation initially suggested different optimal hyper-parameters for BCE and Poisson loss (lr = 0.002, Table 6), these two losses were also tested with optimal hyper-parameters retained with DeepMaxent’s loss on PA data to ensure fairness. Interestingly, the results confirmed that the best-performing hyper-parameters were the same as those identified for DeepMaxent’s loss (Table 7), suggesting that these settings were optimal across different loss functions. This raises an important question about how to properly perform cross-validation when working with presence-only (PO) data, which may contain biases due to sampling effort or environmental factors. Since models are ultimately evaluated using AUC on unbiased presence-absence (PA) test data, there is a risk that hyper-parameters optimised on biased PO data may not generalize well to true PA conditions. Ideally, cross-validation should be performed using PA data, at least in part, before applying the models to the PA test set. This would help ensure that the chosen hyper-parameters lead to models that generalize well to unbiased evaluation data, reducing the risk of overfitting to PO-specific biases.
Appendix D Sensitivity analyses
D.1 Sensitivity to the mini-batch size
The impact of mini-batch size on the average AUC values obtained by the DeepMaxent method in the different regions of the dataset is illustrated in Table 8.
Table 8: Impact of mini-batch size on DeepMaxent performance across different regions. The table presents the accuracy scores for each region (AWT, CAN, NSW, NZ, SA, SWI) and the average (avg) performance across all regions.
Mini-batch size regions